OpenStudy (anonymous):

Simplify : 325/x - 325/x+15

TheSmartOne (thesmartone):

simplify 325/x - 325/x = ??

OpenStudy (anonymous):

yes ! I know that I need to find the LCD of x and x+15 but I dont remember how ...

TheSmartOne (thesmartone):

oh you just multiply one fraction both numerator and denominator by x, and the other by (x+15)

OpenStudy (anonymous):

Could you please show me ?

OpenStudy (anonymous):

i would say they end up looking like this : 340/ x+15 - 325/ x+15

TheSmartOne (thesmartone):

\(\huge\frac{325 (x+15)}{x(x+15)} - \frac{325(x)}{(x+15)(x)}=?\)

OpenStudy (anonymous):

Okay so what is he LCD ?

TheSmartOne (thesmartone):

x(x+15)